OpenStudy (tabithax3):

Determine whether the function is even, odd, or neither: f(x)=x^3 + cosx

hartnn (hartnn):

put x = -x in f(x)

hartnn (hartnn):

if f(-x)=f(x) then the function is even

OpenStudy (anonymous):

even

hartnn (hartnn):

if f(-x)=-f(x) then the function is odd

OpenStudy (tabithax3):

@hartnn @ilikephysics2 the back of the book says neither. i don't understand how they got the answer

hartnn (hartnn):

did u put x= -x and neither is correct.

hartnn (hartnn):

what is (-x)^3

OpenStudy (tabithax3):

-x

hartnn (hartnn):

(-x)^3=-x^3 and whats cos(-x) ?

OpenStudy (tabithax3):

cos(-x)=cosx?

hartnn (hartnn):

yup, so f(-x)= -x^3+cos x now is this f(x) ?? is this -f(x) ??

OpenStudy (anonymous):

oh its cause it crosses the origin with respect to x

OpenStudy (anonymous):

do you have a calculator near you?

OpenStudy (tabithax3):

@ilikephysics2 i have both a scientific and graphing calculator with me

OpenStudy (anonymous):

graph it and you will see how it is neither it goes throw the x axis and crosses with respect to y

hartnn (hartnn):

calculator not required, u got f(-x)= -x^3+cos x u have f(x)=x^3+cos x and -f(x)= -x^3-cos x right? so any of these are equal ??

OpenStudy (tabithax3):

@hartnn i don't think any of them are equal, but i might be wrong?

hartnn (hartnn):

you are correct, since \(f(-x) \ne f(x)\) and \(f(-x) \ne -f(x)\) so its neither odd nor even

OpenStudy (tabithax3):

@hartnn thank you for helping, i understand it now :) my professor never explained it in class and my textbook doesn't explain it as well

hartnn (hartnn):

glad to hear that u understand :) welcome.